# Exact Fisher zeros and thermofield dynamics across a quantum critical point

### v2

@@ Abstract
By setting the inverse temperature $\beta$ loose to occupy the complex plane, Michael E. Fisher showed that the zeros of the complex partition function $Z$, if approaching the real $\beta$ axis, reveal a thermodynamic phase transition. More recently, Fisher zeros have been used to mark the dynamical phase transition in quench dynamics. The success of Fisher zeros however seems limited, and it is unclear how they can be employed to shed light on quantum phase transitions or the non-unitary dynamics of open quantum systems. Here we answer this question by a comprehensive analysis of the (analytically continued) one-dimensional transverse field Ising model. We exhaust all the Fisher zeros to show that in the thermodynamic limit they congregate into a remarkably simple pattern in the form of continuous open or closed lines. These Fisher lines evolve smoothly as the coupling constant is tuned, and a qualitative change identifies the quantum critical point. By exploiting the connection between $Z$ and the thermofield double states, we obtain analytical expressions for the short- and long-time dynamics of the survival amplitude and the scaling of recurrence time at the quantum critical point. We further point out $Z$ can be realized and probed in monitored quantum circuits. The analytical results are corroborated by numerical tensor renormalization group which elevates the approach outlined here to a powerful tool for interacting quantum systems.
@@
There is a renewed interest in the concept of complex-valued partition function and its zeros in quantum many-body physics. In the seminal work of Lee and Yang on phase transitions, the magnetic field in spin Hamiltonians (or the chemical potential in a grand canonical ensemble) is analytically continued to take on complex values [1, 2]. Then phase transitions in the thermodynamic limit can be inferred and analyzed by tracking the zeros of the complex-valued partition function $Z$ evaluated for finite-size systems. Fisher extended the recipe by allowing the inverse temperature $\beta=1/k_{B}T$ to be complex instead, which applies to all systems [3]. Yet, compared to the Lee-Yang zeros, the Fisher zeros are more challenging to find or analyze. Because $\beta$ couples to every term in the Hamiltonian, the locations of Fisher zeros on the complex $\beta$ plane appear to feature less regularity. They may not form simple smooth curves as the thermodynamic limit is approached, in sharp contrast to the celebrated circle theorem for Lee-Yang zeros [1, 2]. Such ostensible complexity may be responsible for the scarcity of exact results on Fisher zeros (for a review, see Ref. [4]).
Fisher zeros were invented to comprehend thermodynamic phase transitions. Recently, several of us [5] proposed to use them to probe quantum phase transitions [6]. Numerical data on the one-dimensional transverse field Ising Model (1DTFIM) reveal that as the coupling constant (the transverse field) $g$ is tuned, the motion of certain Fisher zeros correlates with the energy gap associated with the domain wall excitations. The analysis of Ref. [5] was limited to $g\leq 1$, so a natural question is what happens on the disordered side $g>1$. Naively, one might expect from the Kramers-Wannier (KW) duality [7] that the zeros at $g>1$ simply mirror those at $g<1$. This conjecture turns out to be false. We shall show that the KW duality breaks down for complex $\beta$. To establish a correct, complete picture of Fisher zeros across the quantum critical point (QCP), it is desired to have a convenient exact formula of $Z$ valid for all $g$ values. Then analytical expressions of $Z$ in various regions of the complex $\beta$ plane can be derived to delineate the intricate interplay of quantum and thermal fluctuations.
@@
Our results are corroborated by exact analytical solution, exact diagonalization, and high order tensor renormalization group (HOTRG) [16]. While these observations are made for 1DTFIM, we conjecture that the regularity of Fisher’s zeros and the rich implications of the complex partition function ${Z}$ generalize to a host of other many-body systems. Since ${Z}$ and its zeros can be reliably computed for relatively small system sizes, the program outlined here can serve as a powerful diagnostic tool for quantum many-body physics.
@@
While 1DTFIM is exactly solvable, the analytical form of its partition function $Z$ is subtle to write down. Usually the model is treated in fermionic representation after the Jordan-Wigner transformation, where one has to confront the two sectors (even and odd) of the total fermion number that involve $L$-dependence [17]. We find it cleaner to stick to the spin representation and map $Z=\mathrm{Tr}e^{-\beta H}$ to that of an anisotropic 2D Ising model, and its Onsager-Kaufman solution [18, 19], by Trotter decomposition along the imaginary time axis $\beta$ which plays the role of another spatial dimension $y$ [20]. The result is
@@
where $\epsilon_{k}=[1+g^{2}-2g\cos(\pi k/L)]^{1/2}$. We refer to Eq. (2) as the Suzuki solution, even though the all important sign$(g_{c}-g)$ did not appear explicitly in Ref. [20] and perhaps is not widely appreciated. Without the sign factor, one would have expected the KW duality [7], i.e. $Z$ is invariant under the transformation $g\rightarrow 1/g$, $\beta\rightarrow g\beta$ so that it is sufficient to analyze the subspace $g<1$. This is only an illusion: the correct formula Eq. (2) shows that $Z$ itself does not possess KW duality and implies the presence of the recently proposed non-invertible symmetry [21, 22]. The distribution of Fisher zeros on the quantum disordered side $g>1$ is qualitatively different from the ordered side $g<1$.
With the form of $Z$ determined, we now sketch how Fisher zeros are located and visualized. To analytically continue $Z$, we simply allow $\beta$ to be complex, $\beta=\beta_{r}+i\beta_{i}$, in Eq. (2). On the complex $\beta$ plane, $Z(\beta)$ is not a polynomial but an entire function of $\beta$. Following Fisher [3], we can factorize $Z$ into $Z(\beta)=e^{h(\beta)}\prod_{j}(\beta-z_{j})$ where $h$ is an entire function, and the multiplicity of $z_{j}$, the $j$-th root of $Z=0$, is assumed to be 1. Define the free energy (up to a factor of $\beta$) as $f=-\ln Z=-h-f_{na}$, where the non-analytical part $f_{na}=\sum_{j}\ln(\beta-z_{j})$ develops singularities at $\beta=z_{j}$ and is completely determined by the Fisher zeros $\{z_{j}\}$. Let $f_{na}=\phi+i\psi$, then real part $\phi=\sum_{i}\ln|\beta-z_{j}|$ can be interpreted as the electric potential due to a set of point charges at $\{z_{j}\}$ [4]. The electric field lines stem out of these Fisher charges and correspond to the contour line of the imaginary part $\psi$ [4]. A branch cut of $f$ is developed at each Fisher zero. Thus the contour plot of the real and imaginary part of $f$ can serve as an intuitive way to locate the Fisher zeros, matching the exact solution (Fig. 2). Note that in Fig. 2, due to the mirror symmetry of Fisher zeros with respect to the $\beta_{i}$ axis, only the Fisher zeros at the first quadrant is displayed. In the thermodynamic limit, if the Fisher zeros come close to each other, the charge density $\rho(\beta)=\sum_{j}\delta(\beta-z_{j})$ may become a continuous function.
@@
Tensor renormalization group with complex $\beta$. We also compute $Z$ using an independent method which serves two purposes. First, it confirms the analytical solution Eq. (2) is correct. Second, it demonstrates a general and precise technique to compute $Z(\beta\in\mathbb{C})$ for interacting quantum spin models for which there is no exact solution. Since $Z$ can be represented as the trace of a product of local tensors, it can be evaluated by tensor network algorithms [23, 24, 25] which have proven to be efficient and accurate for spin models [26]. Starting from the 1D system, we first construct a 2D “lattice” of size $L\times N$ by Trotter decomposition $\beta=\tau N$. Then, $Z$ is obtained from HOTRG. The results are in agreement with both exact diagonalization and the exact solution above.
@@
It is worth to mention that when $\beta$ is complex, Monte Carlo develops sign problems but HOTRG can yield $Z$ accurately [26]. Previously renormalization group (RG) transformation has been generalized to the complex $\beta$ plane to study confinement in lattice gauge theory and classical $O(N)$ models. It was observed that the Fisher zeros are located at the boundary of the attraction basins of infrared fixed points, and they control the global properties of the complex RG flows [27, 28]. Here we confine our attention to the evaluation of Fisher zeros through HOTRG but for a quantum spin model.
Figure 2: Locating Fisher zeros. (a) The contour lines of ${\rm Re}f$. The white lines indicate the branch cuts, which terminate the Fisher zeros at the starting points ($L=32$,$g=1.8$).(b)Zeros of the same system determined by the intersections of ${\rm Re}(Z)=0$ and ${\rm Im}(Z)=0$.
@@
Motion of Fisher zeros across QCP. Fig. 1 sketches the evolution of Fisher zeros on the complex $\beta$ plane as the coupling constant $g$ is varied. The main observations are summarized into (a) to (d) below. (a) The Fisher zeros of 1DTFIM appear as continuous lines, rather than isolated points or densely packed regions, in the thermodynamic limit $L\rightarrow\infty$. (b) For $g<1$, there exist open lines of zeros that are approximately parallel to the $\beta_{r}$ axis. (c) As $g$ increases from 0 to 1, these open lines shift upwards toward large $\beta_{i}$, and vanish completely at the QCP $g=1$. (d) Away from $g=0$, closed loops of Fisher zeros also appear and persist for all $g$. These loops are centered around the $\beta_{i}$ axis and move towards $\beta_{i}=0$ as $g$ is increased. We stress that properties (a) to (d) came as surprises, in the sense that they are not obvious from staring at the model or the exact solution of $Z$.
@@
Some of these features can be appreciated by considering a few limit cases. At $g=0$, $Z=\mathrm{Tr}V^{L}$ where the transfer matrix $V=e^{\beta}+\sigma_{x}e^{-\beta}$. The Fisher zeros are determined by the condition $\tanh\beta=e^{in\pi/L}$ with $n$ an integer, i.e. they are uniformly distributed on the unit circle on the plane of $\tanh\beta$ [13]. On the complex $\beta$ plane, the zeros form a set of straight lines at $\beta_{i}=\pi/4+m\pi$, with $m$ an integer, as $L\rightarrow\infty$. Another simple case is the small size system with $L=2$. After some algebra, one finds the zeros are located on the $\beta_{i}$ axis at
@@
where $m$ again is an integer. While the size $L=2$ is too small to feature Fisher zeros away from the $\beta_{i}$ axis or the closed loops, it does correctly capture the trend that the zeros move toward $\beta_{i}=0$ as $g$ is increased.
It is clear from (d) that the QCP is marked by a qualitative change in the Fisher zeros. To describe the lowest open Fisher line, one can take its imaginary part $\beta_{i}$ at some fixed large value of $\beta_{r}$ and plot $1/\beta_{i}$ against $g$. In Ref. [5], it was shown that the resulting curve is linear near QCP. Similarly, for the closed Fisher lines, we can select a characteristic point on each curve, e.g. recording the $\beta_{i}$ value of the rightmost point. Then a linear relationship between $1/\beta_{i}$ and $g$ is observed on the quantum disordered side $g>1$, which is consistent with the spin flip excitations. A more refined description is to define the velocities of the zeros on the complex $\beta$ plane with increasing $g$. Interestingly, the speed of the zeros exhibits properties similar to the Grüneisen parameter [29, 30].
In the next section, we will describe another way to determine the QCP from the asymptotic oscillations of $Z$ at low temperatures, which gives consistent results. We conjecture that in general the distribution of Fisher zeros must undergo a qualitative change at QCP. Otherwise, the two quantum phases separated by the QCP would share the same global patterns of RG flows on the complex $\beta$ plane, which would contradict to the fact that the two phases correspond to different RG fixed points.
What can $Z$ tell us about quantum dynamics? We now switch perspective to view $Z$ as the survival amplitude of certain quantum state. Let $\beta=\beta_{r}+it$, and interpret the imaginary (real) part of $\beta$ as time $t$ (the inverse temperature $\beta_{r}=1/k_{B}T$). Let $E_{n}$ be the eigen energies, and $|n\rangle$ the corresponding eigen states, of Hamiltonian $H$. Then
@@
In contrast to the Loschmidt echo in quench dynamics [9], here the connection between $Z$ and quantum dynamics is established by the TFD states, which are dual to the eternal black hole in the context of AdS/CFT correspondence [15]. Via this bridge, the detailed knowledge about the analytical structure of $Z$, including the Fisher zeros, can now be translated into new insights about quantum dynamics. Three examples are given below.
@@
Figure 3: The rich behaviors of the spectral form factor (survival probability) $S(\beta_{r},t)$. (a) The short-time decay of $S$ is dictated by the specific heat capacity $C_{V}/L$. (b) The low temperature oscillations (inset, $g=1.5$) of $S$. The oscillation frequency obtained from Fourier transform can be fit by $\Delta=2(g-1)$ and vanishes at QCP. (c) Fine structures of $S$ over long time scales for different $\beta_{r}$. At the QCP, all exhibit the same periodicity $t^{*}=4L$, the quantum recurrence time.
@@
where the oscillation frequency $\Delta=\sum_{k}(\epsilon_{2k-1}-\epsilon_{2k})$ has the meaning of the excitation gap. Previously in Ref. [5], it was noted from numerics that for $g\leq 1$ the oscillation frequency vanishes at the QCP while the oscillation amplitude reaches maximum. The analytic asymptote derived here not only explains this observation but also expands its validity to the disordered side $g\geq 1$ [Fig. 3(b)]. It seems quite general that the low temperature oscillations of $Z(\beta_{r},t)$ can be used to identify QCP.
The dynamics at the QCP exhibits an interesting scaling behavior and a rather rich fine structure. For a system of finite size $L$, the energy spectrum is discrete and the initial TFD state will come back to itself after the quantum recurrence time $t^{*}$. Previously, it was claimed that $t^{*}=8L$ [17]. Here our numerical and analytical calculations consistently predict $t^{*}=4L$. For example, take the limit of large $\beta_{r}$, expand the exact Suzuki solution and carry out the $k$-sum, we find at $g=1$ the oscillation frequency is $\Delta=2\tan({\pi}/{4L})$ which becomes ${\pi}/{2L}$ in the limit of large $L$ to yield $t^{*}=4L$. As shown in Fig. 3(c), for smaller $\beta_{r}$, $S(t)$ deviates significantly from the sinusoidal form to acquire fine structures, but $t^{*}$ stays approximately the same. We also observe that at the QCP and for intermediate or large $(\beta_{r},t)$, $S$ exhibits self-similarity, i.e. $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ with $n$ an integer. This is consistent with the notion of criticality and provides a nice way to define RG flow in the complex plane (see Supplementary Materials).
@@
In summary, this work presents a comprehensive analysis of the Fisher zeros in the 1DTFIM. We have demonstrated that Fisher zeros form continuous, smoothly shifting curves in the complex $\beta$ plane, and they offer insights into quantum criticality despite never touching the real $\beta$ axis. The complex partition function $Z$ provides valuable analytical predictions about the quantum dynamics which is beyond the reach of the traditional thermodynamics, and can be realized and probed in nonunitary quantum circuits. Our improvement/clarification over the original Suzuki solution correctly captures the different excitations on the two sides of the QPT and the breakdown of the Kramers-Wannier duality. These findings demonstrate the power of this approach, and open up a new avenue to study quantum critical systems in 1D [41, 42, 43] and higher dimensions [44, 45, 46].

### v3

@@ Abstract
By setting the inverse temperature $\beta$ loose to occupy the complex plane, Michael E. Fisher showed that the zeros of the complex partition function $Z$, if approaching the real $\beta$ axis, reveal a thermodynamic phase transition. More recently, Fisher zeros were used to mark the dynamical phase transition in quench dynamics. It remains unclear, however, how Fisher zeros can be employed to better understand quantum phase transitions or the non-unitary dynamics of open quantum systems. Here we answer this question by a comprehensive analysis of the analytically continued one-dimensional transverse field Ising model. We exhaust all the Fisher zeros to show that in the thermodynamic limit they congregate into a remarkably simple pattern in the form of continuous open or closed lines. These Fisher lines evolve smoothly as the coupling constant is tuned, and a qualitative change identifies the quantum critical point. By exploiting the connection between $Z$ and the thermofield double states, we obtain analytical expressions for the short- and long-time dynamics of the survival amplitude including its scaling behavior at the quantum critical point. We point out $Z$ can be realized and probed in monitored quantum circuits. The exact analytical results are corroborated by numerical tensor renormalization group. We further show similar patterns of Fisher zeros also emerge in other spin models. Therefore the approach outlined may serve as a powerful tool for interacting quantum systems.
@@
There is a renewed interest in the concept of complex-valued partition function and its zeros in quantum many-body physics. In the seminal work of Lee and Yang on phase transitions, the magnetic field in spin Hamiltonians (or the chemical potential in a grand canonical ensemble) is analytically continued to take on complex values [1, 2]. Then phase transitions in the thermodynamic limit can be inferred and analyzed by tracking the zeros of the complex-valued partition function $Z$ evaluated for finite-size systems. Fisher extended the recipe by allowing the inverse temperature $\beta=1/k_{B}T$ to be complex instead, which applies to all systems [3]. Yet, compared to the Lee-Yang zeros, the Fisher zeros are more challenging to find or analyze. Because $\beta$ couples to every term in the Hamiltonian, the locations of Fisher zeros on the complex $\beta$ plane appear to feature less regularity. They may not form simple smooth curves as the thermodynamic limit is approached [4], in sharp contrast to the celebrated circle theorem for Lee-Yang zeros [1, 2]. To unveil the unifying patterns behind the ostensible complexity of Fisher zeros, exact analytical results would be greatly helpful.
@@
Fisher zeros were invented to comprehend thermodynamic phase transitions. Recently, several of us [5] proposed to use them to probe quantum phase transitions [6]. Numerical data on the one-dimensional transverse field Ising Model (1DTFIM) reveal that as the coupling constant (the transverse field) $g$ is tuned, the motion of certain Fisher zeros correlates with the energy gap associated with the domain wall excitations. The analysis of Ref. [5] was limited to $g\leq 1$, so a natural question is what happens on the disordered side $g>1$. Naively, one might expect from the Kramers-Wannier (KW) duality [7] that the zeros at $g>1$ simply mirror those at $g<1$. This conjecture turns out to be false. We shall show that the KW duality breaks down for complex $\beta$. To establish a correct, complete picture of Fisher zeros across the quantum critical point (QCP), an exact formula of $Z$ valid for all $g$ values is required to delineate the intricate interplay of quantum and thermal fluctuations in various regions of the complex $\beta$ plane.
@@
While 1DTFIM is exactly solvable, the analytical form of its partition function $Z$ is subtle to write down. Usually the model is treated in fermionic representation after the Jordan-Wigner transformation, where one has to confront the two sectors of the total fermion number [16]. We find it cleaner to stick to the spin representation and map $Z=\mathrm{Tr}e^{-\beta H}$ to that of an anisotropic 2D Ising model and its Onsager-Kaufman solution [17, 18], by Trotter decomposition along the imaginary time axis $\beta$ which plays the role of another spatial dimension $y$ [19]. The result is
@@
where $\epsilon_{k}=[1+g^{2}-2g\cos(\pi k/L)]^{1/2}$. We refer to Eq. (2) as the Suzuki solution, even though the all important sign$(g_{c}-g)$ did not appear explicitly in Ref. [19] and perhaps is not widely appreciated. Without the sign factor, one would have expected the KW duality [7], i.e. $Z$ is invariant under the transformation $g\rightarrow 1/g$, $\beta\rightarrow g\beta$ so that it is sufficient to analyze the subspace $g<1$. This is only an illusion: the correct formula Eq. (2) shows that $Z$ itself does not possess KW duality and implies the presence of the recently proposed non-invertible symmetry [20, 21]. The distribution of Fisher zeros on the quantum disordered side $g>1$ is qualitatively different from the ordered side $g<1$. From the expression of $Z$, Fisher zeros can be located and visualized by following the procedure outlined in [22].
@@
Tensor renormalization group with complex $\beta$. We also compute $Z$ using an independent numerical method which serves two purposes. First, it confirms the analytical solution Eq. (2) is correct. Second, it demonstrates a general and precise technique to compute $Z(\beta\in\mathbb{C})$ for interacting quantum spin models for which there is no exact solution. Since $Z$ can be represented as the trace of a product of local tensors, it can be evaluated by tensor network algorithms [23, 24, 25] which have proven to be efficient and accurate for spin models [26]. Starting from the 1D system, we first construct a 2D “lattice” of size $L\times N$ by Trotter decomposition $\beta=\tau N$. Then, $Z$ is obtained from High Order Tensor Renormalization Group (HOTRG) [27]. We have checked that tensor results are in agreement with both exact diagonalization and the exact solution above (see [22] for details).
@@
It is worth to mention that when $\beta$ is complex, Monte Carlo develops sign problems but HOTRG can yield $Z$ accurately [26]. Previously renormalization group (RG) transformation has been generalized to the complex $\beta$ plane to study confinement in lattice gauge theory and classical $O(N)$ models. It was observed that the Fisher zeros are located at the boundary of the attraction basins of infrared fixed points, and they control the global properties of the complex RG flows [28, 29]. Here we confine our attention to the evaluation of Fisher zeros through HOTRG but for quantum spin models.
@@
Motion of Fisher zeros across QCP. The exact solution corroborated by HOTRG yields a complete picture, sketched in Fig. 1, of the evolution of Fisher zeros on the complex $\beta$ plane as the coupling constant $g$ is varied. The main observations are summarized into (a) to (d) below. (a) The Fisher zeros of 1DTFIM appear as continuous lines, rather than isolated points or densely packed regions, in the thermodynamic limit $L\rightarrow\infty$. (b) For $g<1$, there exist open lines of zeros that are approximately parallel to the $\beta_{r}$ axis. (c) As $g$ increases from 0 to 1, these open lines shift upwards toward large $\beta_{i}$, and vanish completely at the QCP $g=1$. (d) Away from $g=0$, closed loops of Fisher zeros also appear and persist for all $g$. These loops are centered around the $\beta_{i}$ axis and move towards $\beta_{i}=0$ as $g$ is increased. We stress that properties (a) to (d) came as surprises, in the sense that they are not obvious from staring at the model or the exact solution of $Z$.
@@
Some of these features can be appreciated by considering a few limit cases. At $g=0$, $Z=\mathrm{Tr}V^{L}$ where the transfer matrix $V=e^{\beta}+\sigma_{x}e^{-\beta}$. The Fisher zeros are determined by the condition $\tanh\beta=e^{in\pi/L}$ with $n$ an integer, i.e. they are uniformly distributed on the unit circle on the plane of $\tanh\beta$ [13]. On the complex $\beta$ plane, the zeros form a set of straight lines at $\beta_{i}=\pi/4+m\pi/2$, with $m$ an integer, as $L\rightarrow\infty$. Another simple case is the small size system with $L=2$. After some algebra, one finds the zeros are located on the $\beta_{i}$ axis at
@@
where $m$ again is an integer. While the size $L=2$ is too small to feature Fisher zeros away from the $\beta_{i}$ axis or the closed loops, it does correctly capture the trend that the zeros move toward $\beta_{i}=0$ as $g$ is increased. We note that previous work [30] focused on zeros on the imaginary $\beta$ axis originated from the $\cosh$ terms in $Z$.
It is clear from (c) that the QCP is marked by a qualitative change in the Fisher zeros. To describe the lowest open Fisher line, one can take its imaginary part $\beta_{i}$ at some fixed large value of $\beta_{r}$ and plot $1/\beta_{i}$ against $g$. In Ref. [5], it was shown that the resulting curve is linear near QCP. Similarly, for the closed Fisher lines, we can select a characteristic point on each curve, e.g. recording the $\beta_{i}$ value of the rightmost point. Then a linear relationship between $1/\beta_{i}$ and $g$ is observed on the quantum disordered side $g>1$, which is consistent with the spin flip excitations. A more refined description is to define the velocities of the zeros on the complex $\beta$ plane with increasing $g$. Interestingly, we find the speed of the zeros exhibits properties similar to the Grüneisen parameter [31, 32], see [22].
Figure 2: The pattern of Fisher zeros, as the intersections of $Z_{r}=0$ (red) and $Z_{i}=0$ (blue) computed by HOTRG for finite-size systems ($L=32$), on the complex $\beta$ plane. (a) The zeros for the ANNNI model with $\kappa=0.1$ and $g=0.5$, calculated with D-bond $D_{b}=30$, form an open line near the $\beta_{r}$-axis. (b) In the XY model with $\gamma=0.01$, calculated with $D_{b}=50$, the zeros near the $\beta_{r}$-axis form closed loops.
Fisher zeros in other spin models. We conjecture that the regularity of Fisher’s zeros and its correlation with QCP, observed above for 1DTFIM, generalize to a host of other systems. Since ${Z}$ and its zeros can be reliably computed from tensor network method, the program outlined here can serve as a powerful diagnostic tool for quantum many-body physics. To illustrate this, we consider two examples. The first is the axial next nearest neighbour Ising (ANNNI) model with the Hamiltonian
$H=-\sum_{i=1}^{L}\left(\sigma_{i}^{z}\sigma_{i+1}^{z}-\kappa\sigma_{i}^{z}\sigma_{i+2}^{z}+g\sigma_{i}^{x}\right),$
where we choose $\kappa=0.1$ and $g=0.5$ so the system has a gap. The second is the $XY$ model with the Hamiltonian
$H=\sum_{i=1}^{L}[(1+\gamma)\sigma_{i}^{x}\sigma_{i+1}^{x}+(1-\gamma)\sigma_{i}^{y}\sigma_{i+1}^{y}],$
where we chose $\gamma=0.01$, bringing the system close to a QCP. For both systems, numerical calculations show that as $L$ increases, the zeros indeed approach continuous lines. Figure 2 compares the Fisher zeros of these two models obtained from HOTRG with $L=32$ [22]. For the ANNNI model, open lines corresponding to domain wall excitations still exist. But the zero locations are no longer symmetric between $\beta_{r}>0$ and $\beta_{r}<0$, causing the open lines to tilt. For the XY model near QCP, we witness the same qualitative change as seen in the 1DTFIM: the open lines completely vanish, giving way to closed loops. The change is mandatory because otherwise, the two phases separated by the QCP would share the same global patterns of RG flows on the complex $\beta$ plane, which would contradict to the fact that they correspond to different RG fixed points.
What can $Z$ tell us about quantum dynamics? We now view $Z$ from another perspective, as the survival amplitude of certain quantum state. Let $\beta=\beta_{r}+it$, and interpret the imaginary (real) part of $\beta$ as time $t$ (the inverse temperature $\beta_{r}=1/k_{B}T$). Let $E_{n}$ be the eigen energies, and $|n\rangle$ the corresponding eigen states, of Hamiltonian $H$. Then
@@
In contrast to the Loschmidt echo in quench dynamics [9], here the connection between $Z$ and quantum dynamics is established by the TFD states, which are dual to the eternal black hole in the context of AdS/CFT correspondence [15]. Via this bridge, the detailed knowledge about the analytical structure of $Z$, including the Fisher zeros, can now be translated into new insights about quantum dynamics. Three examples for the 1DTFIM are given below.
Figure 3: The rich behaviors of the spectral form factor (survival probability) $S(\beta_{r},t)$. (a) The short-time decay of $S$ is dictated by the specific heat capacity $C_{V}/L$. (b) The low temperature oscillations (inset, $g=1.5$) of $S$. The oscillation frequency obtained from Fourier transform can be fit by $\Delta=2(g-1)$ and vanishes at QCP. (c) Fine structures of $S$ over long time scales for different $\beta_{r}$ and $L$. At the QCP, all exhibit the same periodicity $t^{*}=4L$. The overlay of curves shows self-similarity, $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ with $n=2$.
@@
where the oscillation frequency $\Delta=\sum_{k}(\epsilon_{2k-1}-\epsilon_{2k})$ has the meaning of the excitation gap. Previously in Ref. [5], it was noted from numerics that for $g\leq 1$ the oscillation frequency vanishes at the QCP while the oscillation amplitude reaches maximum. The analytic asymptote derived here not only explains this observation but also expands its validity to the disordered side $g\geq 1$ [Fig. 3(b)]. Thus, the low temperature oscillations of $Z(\beta_{r},t)$ provide another diagnostic tool for QCP.
The dynamics of $S$ exhibits rich scaling behaviors at the QCP. For a system of finite size $L$, the energy spectrum is discrete and the initial TFD state will come back to itself after the quantum recurrence time $t^{*}$. To find $t^{*}$, take the limit of large $\beta_{r}$, expand the exact Suzuki solution and carry out the $k$-sum, we find at $g=1$ the oscillation frequency is $\Delta=2\tan({\pi}/{4L})$ which becomes ${\pi}/{2L}$ in the limit of large $L$ to yield $t^{*}=4L$. For smaller $\beta_{r}$, $S(t)$ deviates significantly from the sinusoidal form to acquire fine structures, but $t^{*}$ stays approximately the same [22]. Our prediction $t^{*}=4L$ agrees with the conformal field theory result from Ref [16] (see [22] for details). We also observe that $S$ exhibits self-similarity, i.e. $S(\beta_{r},t)_{L}\approx S(n\beta_{r},nt)_{nL}$ with $n$ an integer, at the QCP and for intermediate to large $(\beta_{r},t)$. Figure. 3(c) shows the example of $n=2$. This is consistent with the notion of criticality and provides a nice way to define RG flow in the complex plane (see [22] for details).
@@
In summary, this work presents a comprehensive analysis of the Fisher zeros in the 1DTFIM. We have demonstrated that Fisher zeros form continuous, smoothly shifting curves in the complex $\beta$ plane, and they offer insights into quantum criticality despite never touching the real $\beta$ axis. The complex partition function $Z$ provides valuable analytical predictions about the quantum dynamics which is beyond the reach of the traditional thermodynamics, and can be realized and probed in nonunitary quantum circuits. Our improvement/clarification over the original Suzuki solution correctly captures the different excitations on the two sides of the QPT and the breakdown of the Kramers-Wannier duality. Successful generalization of the numerical analysis to other (the ANNNI and XY) spin models demonstrates that this approach can open up a promising avenue to study quantum critical systems in 1D [43, 44, 45] and higher dimensions [46, 47, 48].